OpenStudy (anonymous):

Evaluate:

OpenStudy (anonymous):

I know how to do it but I am not sure because it is a negative so would I still do it the same way?

OpenStudy (gorv):

\[\sum_{2}^{10} (-2n) - \sum_{2}^{10} 3\]

OpenStudy (gorv):

\[-\sum_{2}^{10} 2n - \sum_{2}^{10} 3\]

OpenStudy (anonymous):

So wait give me a second to try it my self I got that it equals -135 :)

OpenStudy (gorv):

\[-2\sum_{2}^{10} n - \sum_{2}^{10} 3\]

OpenStudy (gorv):

-2(2+3+4.....+10) - (3+3+3......+3....8 times)

OpenStudy (gorv):

-2*54-27=-108-27=-135

OpenStudy (anonymous):

Yay I got it right thank you sooooooooooo much you were awesome :)